If two dice are rolled one time. Find the probability of getting the results. (A) A sum of 6 ( b) doubles ( c) A sum of 7 or 11 ( d) A sum greater than 9 (e) A sum less or equal.


1
Expert's answer
2021-08-31T12:48:11-0400

Two dice are rolled one time


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c:c}
 & 1 & 2 & 3 & 4 & 5 & 6 \\ \hline
 1 & 1,1 & 1,2 & 1,3 & 1,4 & 1,5 & 1,6 \\
 \hdashline
 2 & 2,1 & 2,2 & 2,3 & 2,4 & 2,5 & 2,6 \\
 \hdashline

 3 & 3,1 & 3,2 & 3,3 & 3,4 & 3,5 & 3,6 \\
 \hdashline
 4 & 4,1 & 4,2 & 4,3 & 4,4 & 4,5 & 4,6 \\
 \hdashline
 5 & 5,1 & 5,2 & 5,3 & 5,4 & 5,5 & 5,6 \\
 \hdashline
 6 & 6,1 & 6,2 & 6,3 & 6,4 & 6,5 & 6,6 \\
 \hline
\end{array}"

There are "6^2=36" outcomes.

a)


"1+5=2+4=3+3=4+2=5+1=5"

"P(sum=6)=\dfrac{5}{36}"

b)


"(1,1), (2,2), (3,3), (4,4), (5,5), (6,6)"

"P(doubles)=\dfrac{6}{36}=\dfrac{1}{6}"

c)


"1+6=2+5=3+4=4+3=5+2=6+1=7"

"5+6=6+5=11"

"P(sum=7\ or\ 11)=\dfrac{6}{36}+\dfrac{2}{36}=\dfrac{2}{9}"

d)


"4+6=5+5=6+4=10>9"

"5+6=6+5=11>9"

"6+6=12>9"

"P(sum>9)=\dfrac{3}{36}+\dfrac{2}{36}+\dfrac{1}{36}=\dfrac{1}{6}"


e)


"1+1=2\leq4"

"1+2=2+1=3\leq4"

"1+3=2+2=3+1=4\leq4"

"P(sum\leq4)=\dfrac{1}{36}+\dfrac{2}{36}+\dfrac{3}{36}=\dfrac{1}{6}"